OpenStudy (anonymous):

solve: log2(x2−x−2)=2

OpenStudy (anonymous):

\[ \log_{2}(x ^{2}-x-2)=2 \]

OpenStudy (campbell_st):

if you raise each as a power of 2 \[2^{\log_{2}(x^2 - x -2)} = 2^2\] you end up with x^2 - x - 2 = 4 since \[a^{\log_{a}(u)} = u\] now you should be able to solve the equation

OpenStudy (anonymous):

x = 3, -2

OpenStudy (anonymous):

thank you guys :)